Directions:Solve each of the equations for x.
Problem# 22
25x^2=13

Directions: Solve each equation for x.
Problem# 32
2(x-5)^2=3
Directions: Solve each equation for x.
Problem#42
x^2+4x+4=7 (Hint:Factor the left-hand side)

Answer by tutorcecilia(2152)

25x^2=13
25x^2/25=13/25
x^2=13/25
sqrt(x^2)=sqrt(13/25)
x=(sqrt13)/5
.
.

2(x-5)^2=3
2(x-5)^2/2=3/2
(x-5)^2=3/2
sqrt(x-5)^2=sqrt(3/2)
x-5=sqrt(3/2)
x-5+5=5+-sqrt(3/2)
.
.
x^2+4x+4=7 (Hint:Factor the left-hand side)
(x+2)(x+2)=7
(x+2)^2=7
sqrt(x+2)^2=sqrt(7)
x+2=sqrt7
x+2-2=-2sqrt(7)
x=-2+-sqrt(7)